In a crystalline solid, atom B gives ccp arrangement and atom A occupy 50% of both voids. (1) find the number of voids and empirical formula of the compound. (2) find the number of formula units if 0.5mol of gives arrangement?

Grade:12

1 Answers

Vikas TU
7 years ago
B = 8*1/8 + 6*1/3 => 1 + 2 => 3
and
A = 0.5*(8 + 4) => 12*1/2 => 6
Empirical formulae => A6B3 => 3(A2B)
Total Number of atoms = > 6 + 3 =9
Then the no. of units = 0.5 * 6 * 10^23/9 => 10^23/3